A $136, 000 trust is to be invested in bonds paying 8%, CDs paying 7%, and mortgages paying 10%. The sum of the investments in bonds and CDs must equal the mortgage investment. To earn $12,040 annual income from the investments, how much should the bank invest in bonds?
Answer by stanbon(75887)

A $136,000 trust is to be invested in bonds paying 8%, CDs paying 7%, and mortgages paying 10%. The sum of the investments in bonds and CDs must equal the mortgage investment. To earn $12,040 annual income from the investments, how much should the bank invest in bonds?
-----
Equations::
B + C + M = 136000
8B+7C+10M = 1204000
B + C - M = 0
--------------------------------------
Use any method you know to find that
B = 48000 (amt for Bonds)
C = 20000
M = 68000
---------------
Cheers,
Stan H.
-----------